I am having major problems with this
5(x-2)^2=3

Answer by stanbon(75887)

5(x-2)^2=3
(x-2)^2 = 3/5
x-2 = sqrt(3/5) or x-2 = -sqrt(3/5)
x = 2 + (1/5)sqrt(15) or x = 2 -(1/5)sqrt(15)
=========================
Cheers,
Stan H.